### v3

@@ 2 A Design-Based Model of an Experiment
Figure 2: Likelihood Varies with the Number of Defiers within the Estimated Fréchet Set in Johnson and Goldstein (2003), Smallest Collection of Likelihoods to Contain 95% of the Mass Excludes 8 and 9 Defiers (Shown in Lighter Shading)
@@
2.3 Proposed Maximum Likelihood Decision Rule
@@
We consider a statistical decision problem in the style of Wald (1949), whereby the decision maker wishes to correctly guess the joint distribution of potential outcomes in the sample, having viewed the outcome of the experiment. In Appendix D.1, we present the formal decision problem. Statistical decision theory provides two benefits in our setting: first, decision theory is natural to apply in our design-based setting, unlike alternative criteria like consistency that depend on large-sample or asymptotic assumptions; second, statistical decision theory provides straightforward methods to quantify the gains from exploiting the variation in our likelihood over other decision rules that ignore such variation.
@@
There are finitely many vectors of four integers that sum to the number of participants in the experiment $n$, so $\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})$ is nonempty. We define the maximum likelihood decision rule as an estimator $\hat{{\boldsymbol{\theta}}}_{MLE}$ that places equal weight on each maximizer of the likelihood:
$\mathbb{P}(\hat{{\boldsymbol{\theta}}}_{MLE}({\boldsymbol{x}})={\boldsymbol{\theta}}\mid{\boldsymbol{x}})=\begin{cases}\frac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})\}\}},&\text{if}\quad{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}}),\\ 0,&\text{otherwise}\end{cases}$
@@ 2.4 Justifications for our Proposed Maximum Likelihood Decision Rule
Our primary motivation for proposing the maximum likelihood decision rule is that maximum likelihood rules are considered “standard” in econometric settings, and maximum likelihood estimators are familiar to practitioners. In the following subsections, we provide three additional justifications: first, that the maximum likelihood decision rule varies systematically with the data; second, that the decision rule is Bayes optimal under the appropriate conditions; and third, that the decision rule maximizes entropy, and is therefore the “least informative” distribution to justify the data.
@@ 2.4.1 The Maximum Likelihood Decision Rule Varies Systematically with the Data
Direct inspection of the likelihood in (2) does not reveal obvious patterns in the maximum likelihood estimate. However, by computing the maximum likelihood estimate for every possible realization of the data in fixed sample sizes, we can observe clear patterns in the estimates. Figure 3 illustrates how the maximum likelihood estimates from our proposed decision rule vary systematically with all possible realizations of the data in completely randomized experiments with 50 and 200 people such that half are in intervention. In the subfigure for each sample size, the horizontal axis gives the takeup count in intervention, and the vertical axis gives the takeup count in control. Each cell represents a possible realization of the data, which specifies an estimated Fréchet set in terms of takeup count in intervention and control. The visualization demonstrates that an estimated Fréchet set can also be specified with the estimated average effect, which increases diagonally downward, and the total takeup count in the sample, which increases diagonally upward. An estimated Fréchet set can also be specified with an estimated average effect and a takeup rate in control, as is common in empirical practice.
@@
The figure shows clear patterns in how the MLEs vary with the data, which are visibly similar when we increase the sample size by a factor of four. MLEs in the corners only include one type: always takers if takeup is full in intervention and control, compliers if takeup is full in intervention and zero in control, defiers if takeup is zero in intervention and full in control, and never takers if takeup is zero in intervention and control. MLEs along the edges only include the two types in the corners that they connect.
@@
In the lens-shaped region around the upward-sloping diagonal, MLEs only include always and never takers: the two types in the null hypothesis of the Fisher (1935) exact test. In samples of 50 and 200, the region in which Fisher’s exact test fails to reject the null hypothesis at the 5% level is depicted within the grey dashed line. It is larger than the region in which the MLE includes only always and never takers in both sample sizes.
@@
The figure illustrates how our design-based maximum likelihood decision rule reveals evidence that is obscured by Fisher’s exact test. Fisher’s exact test only considers the null hypothesis that the joint distribution of potential outcomes includes only always and never takers. If it rejects, it does not suggest an alternative. If it fails to reject, a joint distribution of potential outcomes that include types other than always and never takers can maximize the likelihood. With our use of statistical decision theory, we can harness the richer evidence conveyed in the figure beyond the grey dashed lines.
@@
In the lens-shaped region around the downward sloping diagonal, MLEs only include compliers and defiers. This region is related to the region in which an exact test of the null hypothesis that everyone is affected does not reject. There are some realizations of the data near the center of the region in which Fisher’s exact test fails to reject the null hypothesis that no one is affected, but our maximum likelihood estimates indicate that everyone is affected. These estimates are relevant to empirical practice for researchers who wonder if they obtained a near-zero estimated average effect not because no one was affected but because everyone was affected.
@@
Sometimes the MLE can be a tie between more than one distribution. MLEs are unique in the sample of 50, but two-way ties occur in the sample of 200. In the cells bisected by the lines indicated with asterisks–the midlines in which takeup is exactly half in intervention or control–ties occur between the MLEs in the cells on either side of the line. Only at the edges of the figure do the ties result in a unique MLE because the MLE is the same on either side of the line.
@@
The MLEs include three types in all other regions, two regions for each combination. Even though the maximizer of the likelihood only includes two types in our stylized example from Figure 1, maximization of the likelihood sometimes requires trading off between the higher likelihood of fewer types and the higher likelihood of more balance between intervention and control within each type. If restricting the sample to two types requires too much imbalance between intervention and control, the maximizer of the likelihood includes three types. [^2] Viewing Pascal’s triangle as a representation of $N$ choose $K$ can help us to visualize the tradeoff between fewer types (moving to the next row of the triangle, increasing $N$ within each type) and more balance within each type (moving to the middle of the triangle within a row, moving $K$ closer to $N/2$ within each type). We thank Elizabeth Ananat for this point. There are no regions in which a distribution that maximizes the likelihood contains all four types. Therefore, many joint distributions of potential outcomes never maximize the likelihood for any realization of the data.
@@
The shading demonstrates that the MLE reveals evidence beyond the average effect: holding the estimated average effect constant along upward sloping diagonal lines, MLEs can include zero or many defiers. Copas (1973) claims that the likelihood is always maximized at a joint distribution of potential outcomes that preserves the direct estimates of the marginal distributions of potential outcomes from equation (4) and which has either the maximal or minimal number of defiers. In practice, we often find that the likelihood is maximized at a distribution that does not preserve these marginal estimates. However, in all applications we have considered, the likelihood is maximized at either the estimated upper or lower Fréchet bound on defiers.
@@ 2.4.2 The Maximum Likelihood Decision Rule is Bayes Optimal
The maximum likelihood decision rule is Bayes optimal under appropriate specifications of the decision maker’s utility function and subjective prior. [^3] We thank Andriy Norets and Thomas Weimann for bringing the standard conditions that establish Bayes optimality to our attention. For the former, we specify a discrete utility function, which assigns zero utility for incorrect guesses and an arbitrary amount of strictly positive utility for correct guesses; for the latter, we specify a prior that is uniform over all possible distributions of potential outcomes for $n$ subjects. Appendix D.2 reviews the formal definition and proof of Bayes optimality under these conditions.
While the conditions under which our decision rule is optimal are potentially restrictive, we emphasize that the design-based likelihood can be used to compute the Bayes optimal decision rule for any given loss function and prior, as we review in Appendix D. To justify our adoption of a uniform prior, we envision the sample as drawn IID from some population with an unknown joint distribution of potential outcomes (that is, the population shares of always takers, compilers, defiers, and never takers). We adopt a Dirichlet prior over the population distribution of potential outcomes, which is a common choice in sampling models, as it is the conjugate prior of the categorical and multinomial distributions. Our uniform prior over the sample distribution of potential outcomes is consistent with a “flat” Dirichlet prior, with all concentration parameters set to one. Sampling inspired priors like this do not, in general, lead to flat posteriors within Fréchet sets. Regardless of the choice of prior, the ratio of the prior beliefs between any two distributions updates as long as the likelihood ratio differs from one. Adoption of a Bayesian perspective also allows us to construct credible sets for the joint distribution of potential outcomes. In Appendix E, we describe our algorithm for constructing smallest credible sets for the joint distribution of potential outcomes, as well as for the number of each type individually.
@@
Bayes optimality ensures that our proposed decision rule does at least as well as all other decision rules under these optimality conditions. To quantify the gains from utilizing the optimal rule, we compare the Bayes expected utility obtained under the optimal rule to two alternative decision rules. The first alternative decision rule, which we refer to as the “Fréchet rule,” estimates the marginal distributions of potential outcomes directly according to (4), and then chooses randomly among the joint distributions of potential outcomes in the corresponding Fréchet set. The Fréchet rule $\hat{{\boldsymbol{\theta}}}_{F}$ satisfies:
@@
$\mathbb{P}(\hat{{\boldsymbol{\theta}}}_{F}({\boldsymbol{x}})={\boldsymbol{\theta}})=\begin{cases}\frac{1}{\#\mathcal{F}(\hat{\theta}_{1\bullet}({\boldsymbol{x}}),\hat{\theta}_{\bullet 1}({\boldsymbol{x}});n)}&\text{if}\;{\boldsymbol{\theta}}\in\mathcal{F}(\hat{\theta}_{1\bullet}({\boldsymbol{x}}),\hat{\theta}_{\bullet 1}({\boldsymbol{x}});n)\\ 0&\text{otherwise}.\end{cases}$
@@
This decision rule captures the behavior of a statistician who assumes that the experimental data is uninformative within Fréchet sets and is therefore indifferent between any two distributions within the same Fréchet set.
The second alternative decision rule, which we refer to as the “monotonicity rule,” chooses the maximizer of the design-based likelihood among those distributions with zero defiers and/or zero compilers: let $\widehat{{\boldsymbol{\Theta}}}_{M}({\boldsymbol{x}})$ represent the set of joint distributions that maximize the likelihood among those distributions with zero compilers and/or defiers:
$\widehat{{\boldsymbol{\Theta}}}_{M}({\boldsymbol{x}})=\arg\max_{\boldsymbol{\theta}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})\;\text{s.t.}\;\theta_{01}=0\ \text{or}\ \theta_{10}=0.$
Then, the monotonicity decision rule $\hat{{\boldsymbol{\theta}}}_{M}$ satisfies:
$\mathbb{P}(\hat{{\boldsymbol{\theta}}}_{M}({\boldsymbol{x}})={\boldsymbol{\theta}})=\begin{cases}\frac{1}{\#\widehat{{\boldsymbol{\Theta}}}_{M}({\boldsymbol{x}})}&\text{if}\;{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{M}({\boldsymbol{x}})\\ 0&\text{otherwise}.\end{cases}$
This decision rule approximates the behavior of a statistician who adopts a within-sample analogue to the monotonicity assumption of Imbens and Angrist (1994). This rule assumes that the statistician still knows the likelihood function, and is thus the “best” decision rule among those that adopt this monotonicity assumption.
@@
We calculate the exact Bayes expected utility achieved by our optimal maximum likelihood rule and each of the two alternative rules. For a given sample size, we enumerate each of the finitely many feasible sample distributions of potential outcomes $\boldsymbol{\theta}$}. For each distribution $\boldsymbol{\theta}$}, we compute the expected utility of our decision rules by comparing the distribution of estimates they produce under each data realization $\boldsymbol{x}$} to the true distribution. We average the resulting utilities over the values of $\boldsymbol{x}$} according to the likelihood. We then average these expected utilities according to a uniform prior for $\boldsymbol{\theta}$} to get the Bayes expected utility (see Appendix D for additional detail).
@@
Figure 4 plots the ratio of the Bayes expected utility achieved by our proposed maximum likelihood rule to the Bayes expected utility achieved by each of the two alternative rules for even sample sizes between 2 and 50. The gains from using the optimal rule over either of the alternatives increase with the sample size. In a sample size of 50, the Bayes expected utility of the proposed maximum likelihood rule is 1.5 times higher than that of the Fréchet rule, and 1.19 times higher than that of the monotonicity rule.

### v4

@@ 2 A Design-Based Model of an Experiment
Figure 2: The Likelihood Varies with the Number of Defiers within the Estimated Fréchet Set in Johnson and Goldstein (2003), And The Smallest Collection of Likelihoods to Contain 95% of the Mass Excludes 9 Defiers (in Lighter Shading)
@@
2.3 The Design-Based Likelihood Enables Our Proposed Maximum Likelihood Decision Rule
@@
We consider a statistical decision problem in the style of wald1949, whereby the decision maker wishes to correctly guess the joint distribution of potential outcomes in the sample, having viewed the outcome of the experiment. In Appendix C.1, we present the formal decision problem. Statistical decision theory provides two benefits in our setting: first, decision theory is natural to apply in our design-based setting, unlike alternative criteria like consistency that depend on large-sample or asymptotic assumptions; second, statistical decision theory provides straightforward methods to quantify the gains from exploiting the variation in our likelihood over other decision rules that ignore such variation.
@@
There are finitely many vectors of four integers that sum to the number of participants in the experiment $n$, so $\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})$ is nonempty. We define the maximum likelihood decision rule as an estimator $\widehat{{\boldsymbol{\theta}}}_{MLE}$ that places equal weight on each maximizer of the likelihood:
$\mathbb{P}(\widehat{{\boldsymbol{\theta}}}_{MLE}({\boldsymbol{x}})={\boldsymbol{\theta}}\mid{\boldsymbol{x}})=\begin{cases}\dfrac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})\}},&\text{if}\quad{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}}),\\ 0,&\text{otherwise,}\end{cases}$
@@ 2.4 Justifications for our Proposed Maximum Likelihood Decision Rule
Our primary motivation for proposing the maximum likelihood decision rule is that maximum likelihood rules are considered “standard” in econometric settings, and maximum likelihood estimators are familiar to practitioners. In the following subsections, we provide three additional justifications: first, that the maximum likelihood decision rule varies systematically with the data; second, that the decision rule is Bayes optimal under the appropriate conditions; and third, that the decision rule maximizes entropy, and therefore “assumes the least” (jaynes1968) about the joint distribution of potential outcomes.
@@ 2.4.1 The Maximum Likelihood Decision Rule Varies Systematically with the Data
Direct inspection of the likelihood in (2) does not reveal obvious patterns in the maximum likelihood estimate. However, by computing the maximum likelihood estimate for every possible realization of the data in fixed sample sizes, we can observe clear patterns in the estimates. Figure 3 illustrates how the maximum likelihood estimates from our proposed decision rule vary systematically with all possible realizations of the data in completely randomized experiments of 50 and 200 subjects with half in intervention. In a sample of 200 with half in intervention, there are $10,201=101\times 101$ possible realizations of the data and $1,373,301={203\choose 3}$ possible joint distributions of potential outcomes, so the visualization reflects results of grid searches over $14,009,043,501=10,201\times 1,373,301$ total elements.
In the subfigure for each sample size, the horizontal axis gives the takeup count in intervention, and the vertical axis gives the takeup count in control. Each cell represents a possible realization of the data, which specifies an estimated Fréchet set in terms of takeup count in intervention and control. The visualization demonstrates that an estimated Fréchet set can also be specified with the estimated average effect, which increases diagonally downward, and the total takeup count in the sample, which increases diagonally upward. An estimated Fréchet set can also be specified with an estimated average effect and a takeup rate in control, as is common in empirical practice.
@@
The figure shows clear patterns in how the MLEs vary with the data, which are visibly similar when we increase the sample size by a factor of four. MLEs in the corners only include one type: always takers if takeup is full in intervention and control, compliers if takeup is full in intervention and zero in control, defiers if takeup is zero in intervention and full in control, and never takers if takeup is zero in intervention and control. MLEs along the edges and diagonals only include the two types in the corners they connect.
@@
In the convex lens-shaped region around the upward-sloping diagonal, MLEs only include always and never takers: the two types in the null hypothesis of the fisher1935 exact test. In samples of 50 and 200, the region in which Fisher’s exact test fails to reject the null hypothesis at the 5% level is depicted within the grey dashed line. It is larger than the region in which the MLE includes only always and never takers in both sample sizes.
@@
The figure illustrates how our design-based maximum likelihood decision rule reveals evidence that is obscured by Fisher’s exact test. Fisher’s exact test only considers the null hypothesis that the joint distribution of potential outcomes includes only always and never takers. If it rejects, it does not suggest an alternative. If it fails to reject, a joint distribution of potential outcomes that includes types other than always and never takers can maximize the likelihood. With our use of statistical decision theory, we can harness the richer evidence conveyed in the figure beyond the grey dashed lines.
@@
In the convex lens-shaped region around the downward sloping diagonal, MLEs only include compliers and defiers. This region is related to the region in which an exact test of the null hypothesis that everyone is affected does not reject. There are some realizations of the data near the center of the region in which Fisher’s exact test fails to reject the null hypothesis that no one is affected, but our maximum likelihood estimates indicate that everyone is affected. These estimates are relevant to empirical practice for researchers who wonder if they obtained a near-zero estimated average effect not because no one was affected but because everyone was affected.
@@
Sometimes, the MLE can be a tie between more than one distribution. MLEs are unique in the sample of 50 with half in intervention, but two-way ties occur in the sample of 200. In cells bisected by the lines indicated with asterisks, the midlines in which takeup is exactly half in intervention or control, ties occur between MLEs in cells on either side of the line. Only at the edges of the figure do the ties result in a unique MLE because the MLE is the same on either side of the line. For experiments with intervention and control arms unequal in size, ties appear in less predictable patterns.
@@
The MLEs include three types in all other regions, with two regions for each combination. Even though the maximizer of the likelihood only includes two types in our stylized example from Figure 1, maximization of the likelihood sometimes requires trading off between the higher likelihood of fewer types and the higher likelihood of more balance between intervention and control within each type. If restricting the sample to two types requires too much imbalance between intervention and control, the maximizer of the likelihood includes three types. [^1] Viewing Pascal’s triangle as a representation of $N$ choose $K$ can help us to visualize the tradeoff between fewer types (moving to the next row of the triangle, increasing $N$ within each type) and more balance within each type (moving to the middle of the triangle within a row, moving $K$ closer to $N/2$ within each type). We thank Elizabeth Ananat for this point. There are no regions in which a distribution that maximizes the likelihood contains all four types. Therefore, many joint distributions of potential outcomes never maximize the likelihood for any realization of the data. For example, if a sample of 50 consisted of 18 always takers, 22 compliers, 8 defiers, and 2 never takers, the most probable realization of the data would be takeup of 20 in intervention and 13 in control, but the corresponding MLE estimate of 26 always takers, 14, compliers, 0 defiers, and 10 never takers would miss the 8 defiers entirely. We have observed MLEs that contain all four types in experiments with intervention and control arms unequal in size, although we have only observed them tied with MLEs that contain three types.
@@
The shading demonstrates that the MLE reveals evidence beyond the average effect: holding the estimated average effect constant along upward sloping diagonal lines, MLEs can include zero or many defiers. copas1973 claims that the likelihood is always maximized at a joint distribution of potential outcomes that preserves the direct estimates of the marginal distributions of potential outcomes from equation (4) and which has either the maximal or minimal number of defiers. In practice, we often find that the likelihood is maximized at a distribution that deviates slightly from these marginal estimates.
@@ 2.4.2 The Maximum Likelihood Decision Rule is Bayes Optimal
Our proposed maximum likelihood decision rule is Bayes optimal under appropriate specifications of the decision maker’s utility function and prior. For the utility function, we specify a discrete utility function that assigns zero utility for incorrect guesses and an arbitrary amount of strictly positive utility for correct guesses; for the prior, we specify a prior that is uniform over all possible distributions of potential outcomes for $n$ subjects. Appendix C.2 reviews the formal definition and proof of Bayes optimality under these conditions. [^2] We thank Andriy Norets and Thomas Wiemann for bringing the standard conditions that establish Bayes optimality to our attention.
As we review in Appendix C.2, Bayes optimal decision rules can be derived for alternative utility functions and priors through the numerical solution of a linear programming problem, but we offer motivations for the conditions we have chosen, acknowledging their subjectivity. For the utility function, one motivation is that we can interpret it as the payoff function in a game that rewards “being correct,” as in the well-known Monty Hall game, where the contestant receives a prize only if they guess the correct door. Under this utility function, the maximum likelihood decision rule is the Maximum A Posteriori (MAP) rule, which chooses the mode of the posterior distribution. Under a uniform prior, the posterior distribution is proportional to the likelihood, and the MAP estimate equals the MLE estimate.
The desirable feature of our uniform prior is that it assigns equal weight to all distributions in the same Fréchet set, so it is “uninformative” in the sense that it does not confer any information about the joint distribution of potential outcomes beyond their marginal distributions. Furthermore, we can view our uniform prior as the result of sampling from an underlying population over which we hold an uninformative hyperprior. In sampling settings, it is common to assume a Dirichlet hyperprior over population shares. We choose a Dirichlet hyperprior over the population shares of always takers, compliers, defiers, and never takers with all concentration parameters set to one, which is equivalent to an uninformative uniform distribution for the population. Under standard assumptions on sampling from the population, the potential outcomes of subjects are independently and identically distributed (IID), and the sample distribution of potential outcomes follows a multinomial law conditional on the population distribution. Combining the Dirichlet hyperprior and the sampling process yields a Dirichlet-multinomial prior over the counts of each type in the sample. The Dirichlet-multinomial prior with concentration parameters set to one is equivalent to our uniform prior. In our stylized example in Figure 1, our uniform prior implies that the prior is the same in all three rows, so the updated posterior is proportional to the likelihood.
Regardless of the choice of prior, the ratio of the prior beliefs between any two distributions updates as long as the likelihood ratio differs from one, including within Fréchet sets. For example, consider an alternative prior that is uniform over each type for each person in the sample. This prior arises from a highly informative degenerate hyperprior that assumes the population shares of each type are equal. The resulting prior over sample distributions varies within the sample Fréchet sets, and is therefore also informative. In our stylized example in Figure 1, the resulting prior is 0.0220 for the first row, 0.0293 for the second row, and 0.0037 for the third row. The updated posterior distribution maintains curvature within the Fréchet set and updates towards the distribution in the third row that contains two defiers. To achieve a flat posterior distribution across the rows depicted in Figure 1, the prior for the first row with no defiers would have to be exactly 0.75 (=(20/8)/(20/6)) times as large as the second row with one defier and 1.5 (=(20/8)/(20/12)) times as large as the prior for the third row with two defiers.
A Bayesian perspective allows us to construct credible sets for the joint distribution of potential outcomes. In Appendix D, we describe our algorithm for constructing smallest credible sets for the joint distribution of potential outcomes. Smallest credible sets are difficult to characterize because they are multidimensional, so we also describe an algorithm for constructing smallest credible sets for the number of each type individually.
@@
Bayes optimality ensures that our proposed decision rule does at least as well as all other decision rules under these optimality conditions. To quantify the gains from utilizing the optimal rule, we compare the Bayes expected utility obtained under the optimal rule to two alternative decision rules. The first alternative decision rule, which we refer to as the “Fréchet rule,” estimates the marginal distributions of potential outcomes directly according to (4), and then chooses randomly among the joint distributions of potential outcomes in the corresponding Fréchet set. The Fréchet rule $\widehat{{\boldsymbol{\theta}}}_{F}$ satisfies:
@@
$\mathbb{P}(\widehat{{\boldsymbol{\theta}}}_{F}({\boldsymbol{x}})={\boldsymbol{\theta}})=\begin{cases}\dfrac{1}{\#\mathcal{F}(\hat{\theta}_{1\bullet}({\boldsymbol{x}}),\hat{\theta}_{\bullet 1}({\boldsymbol{x}});n)}&\text{if}\;{\boldsymbol{\theta}}\in\mathcal{F}(\hat{\theta}_{1\bullet}({\boldsymbol{x}}),\hat{\theta}_{\bullet 1}({\boldsymbol{x}});n),\\ 0&\text{otherwise}.\end{cases}$
@@
This decision rule captures the behavior of a statistician who assumes that the experimental data are uninformative within Fréchet sets and is therefore indifferent between any two distributions within the same Fréchet set.
The second alternative decision rule, which we refer to as the “monotonicity rule,” approximates the behavior of a statistician who uses the monotonicity assumption of imbens1994 in the population to estimate the share of compliers or defiers in the sample. The rule uses the observed difference in takeup rate between the intervention and control arms $\Delta$ to approximate the share of compliers or defiers. In particular, if the difference is non-negative, the rule imposes a no-defiers monotonicity restriction, and if the difference is non-positive, it imposes a no-compliers monotonicity restriction.
Under this decision rule, the inferred number of compliers or defiers is given by
$\hat{\theta}_{10}=n\cdot\max\{\Delta,0\},\quad\hat{\theta}_{01}=n\cdot\max\{-\Delta,0\}$
rounded to the nearest whole number. Note that when intervention and control arms are the same size, these expressions are always whole numbers. The remaining types, always takers and never takers, are inferred by assuming balance of type shares across both arms. In the case where there are compliers and no defiers, every treated individual in control must be an always taker, and because assignment is randomized, the proportion of always takers in control should be equal to the proportion of always takers in the sample, giving the estimate $\hat{\theta}_{11}=nx_{C1}/(x_{C1}+x_{C0})$. The remaining individuals must all be never takers. Conversely, if there are defiers and no compliers, every treated individual in intervention must be an always taker, so the estimate of always takers is $\hat{\theta}_{11}=nx_{I1}/(x_{I1}+x_{I0})$ and the remaining individuals are never takers. In summary, the monotonicity rule $\widehat{{\boldsymbol{\theta}}}_{M}$ chooses
$\widehat{{\boldsymbol{\theta}}}_{M}({\boldsymbol{x}})$ $=\big(\hat{\theta}_{11},\hat{\theta}_{10},\hat{\theta}_{01},\hat{\theta}_{00}\big)$
$=\bigg(\textbf{1}_{\{\Delta>0\}}\left[n\frac{x_{C1}}{x_{C1}+x_{C0}}\right]+\textbf{1}_{\{\Delta<0\}}\left[n\frac{x_{I1}}{x_{I1}+x_{I0}}\right],\ n\cdot\max\{\Delta,0\},\ n\cdot\max\{-\Delta,0\},$
$\qquad n-\hat{\theta}_{11}-\hat{\theta}_{01}-\hat{\theta}_{10}\bigg)$
with probability one.
@@
We calculate the exact Bayes expected utility achieved by our optimal maximum likelihood rule and each of the two alternative rules. For a given sample size, we enumerate each of the finitely many feasible sample distributions of potential outcomes $\boldsymbol{\theta}$}. For each distribution $\boldsymbol{\theta}$}, we compute the expected utility of our decision rules by comparing the distribution of estimates they produce under each data realization $\boldsymbol{x}$} to the true distribution. We average the resulting utilities over the values of $\boldsymbol{x}$} according to the likelihood. We then average these expected utilities according to a uniform prior for $\boldsymbol{\theta}$} to get the Bayes expected utility (see Appendix C for additional detail).
@@
Figure 4 plots the ratio of the Bayes expected utility achieved by our proposed maximum likelihood rule to the Bayes expected utility achieved by each of the two alternative rules for even sample sizes between 2 and 200. The gains from using the optimal rule over either of the alternatives increase with the sample size. In a sample size of 200, the Bayes expected utility of the proposed maximum likelihood rule is 1.79 times higher than that of the Fréchet rule, and 1.56 times higher than that of the monotonicity rule.